(-6*x^-2*y)^-2
(9x^6)^3/2
log( base 7) (square root of 7)
log(.1)

Answer by stanbon(57224)

(-6*x^-2*y)^-2
= (-6y/x^2)^-2
= (36y^2/x^4)^-1
= 1/(36y^2/x^4)
= x^4/36y^2
=====================
(9x^6)^3/2
= (3x^3)^3
= 27x^9
=================
log( base 7) (square root of 7)
= (1/2) log(base 7) 7
= (1/2) 1
= 1/2
------------
log(.1)
= log (1/10)
= log1 - log10
= 0 - 1
= -1
==============
Cheers,
Stan H.